Given a function described as equation y = 3x + 4, what is y when x is 1, 2, and 3?


1
Expert's answer
2021-05-07T10:23:04-0400

Consider the equation "y=3x+4"

Require to find "y" when "x=1,2,3"

Now "x=1\Rightarrow y=3(1)+4=7"

"x=2\Rightarrow y=3(2)+4=10"

"x=3\Rightarrow y=3(3)+4=13"

Therefore, the values of "y" are: "y=7,10,13"